# Factuality of Large Language Models in the Year 2024

### v2

@@ 6 Challenges and Future Directions
Language models learn the distribution of a language, not facts.
The training objective of language modeling is to maximize the probability of a sentence, as opposed to that of a factual statement. While capable of generating seemingly coherent and fluent outputs upon convergence, models are not guaranteed to always return factually accurate contents.
Automatic evaluation of factual accuracy of open-ended generations remains challenging.
Existing studies on factuality enhancement use different benchmarks and evaluation metrics, making fair comparisons difficult, motivating the need for a unified automated evaluation framework utilizing the same collection of datasets and metrics.
This can be further decomposed into two questions: (1) how to select a set of datasets that can effectively evaluate LLMs factuality? and (2) how to measure the percentage of factual claims given a model response with minimal time expenditure? Current approaches rely on either human evaluation or results of automated fact-checkers, e,g., FactScore and FacTool Min and et al. (2023); Chern et al. (2023). However, automatically quantifying the quality of automated fact-checkers is itself an open question, resulting in a chicken and egg situation.
@@
Retrievers serve as the core component in RAG systems, and the effectiveness of RAG is largely influenced by the quality (coverage and relevance) of the retrieved documents. Latency and difficulties in gathering the most pertinent evidence are the primary challenges in retrieval. While this is partly due to the inability of ranking algorithms to retrieve such documents, certain facts require information gathered from various sources and multi-hop reasoning.
@@
Mitigation in inference: We observe that models can often generate the correct answer in multiple trials even if some of those attempts are incorrect Tian et al. (2023). This motivates us to ask how to provide an anchor that can guide LLM decoding to the factually correct path?
Iteratively detecting, correcting and generating during generation has been demonstrated to be effective to mitigating hallucinations. If simply correcting the first one or two sentences, how much improvements can we expect for subsequent generations? Can factually correct and relevant sentences, phrases or concepts serve as the anchors?
Timely detection of factual errors: Another interesting option to explore is to only detect falsehoods without consideration of correction, allowing LLMs to incorporate uncertainty information and the contrastive probability leveraged in decoding strategies (e.g., CAD and DoLa) for timely detections without evidence.
Improving efficiency and accuracy of automated fact-checkers: The key breakthrough in effectively evaluating the factual accuracy of LLMs lies in establishing accurate and efficient fact-checkers, in which how to evaluate the quality of retrieved evidence in open search space is of significant importance. Current methods rely on the accuracy of claim verification to deduce the quality of evidence by prompting LLMs. Not only is this computationally expensive, it also tends to be highly sensitive to minor prompt changes and LLM updates (e.g., the same prompt may result in inconsistent outputs). Instead, a small task-specific and well fine-tuned NLI model presents a more viable, robust and cost-efficient option.
@@ 7 Conclusion
We presented an overview on the factuality of LLMs, surveying a number of studies covering topics such as evaluation in Section 3 and improvement methods (applicable at various stages: pre-training, SFT, inference and post-processing) in Section 4, along with their respective challenges. We also identified three major issues and additionally pointed out promising future research directions in Section 6.

### v3

@@ 6 Challenges and Future Directions
Language models learn a language distribution, not facts.
The training objective of language modeling is to maximize the probability of a sentence, as opposed to that of a factual statement. While capable of generating seemingly coherent and fluent outputs upon convergence, models are not guaranteed to always return a factual response.
Automatic evaluation of the factual accuracy of open-ended generations remains challenging.
Existing studies on factuality enhancement use different benchmarks and evaluation measures, making fair comparisons difficult, which motivates the need for a unified automated evaluation framework that uses the same collection of datasets and metrics. Current approaches rely on either human evaluation or results of automated fact-checkers such as FactScore and FacTool Min and et al. (2023); Chern et al. (2023). However, automatically quantifying the quality of automated fact-checkers is itself an open question, resulting in a chicken and egg situation.
@@
Retrievers serve as the core component in RAG systems, and the effectiveness of RAGs is largely influenced by the quality (coverage and relevance) of the retrieved documents. Latency and difficulties in gathering the most pertinent evidence are the primary challenges in retrieval. While this is partly due to the inability of ranking algorithms to retrieve such documents, certain facts require information gathered from various sources and multi-hop reasoning.
@@
Mitigation in inference: We observe that models can often generate a correct answer in multiple trials even if some attempts are wrong Tian et al. (2023). This motivates us to ask how to provide an anchor that can guide LLM decoding to the factually correct path?
Iteratively detecting, correcting, and generating during generation has been demonstrated to be effective to mitigate hallucinations. If simply correcting the first one or two sentences, how much improvements can we expect for subsequent generations? Can factually correct and relevant sentences, phrases or concepts serve as anchors?
Development of better retrieval algorithms: Integrating Retrieval-Augmented Generation (RAG) into Large Language Models (LLMs) is challenging due to the prevalence of unreliable information, such as fake news, on the internet. This compromises the accuracy of the knowledge retrieved, resulting in LLMs generating responses based on incorrect input. Consequently, future research should focus on improving retrieval techniques to enhance the factuality of LLM-generated responses.
Improving the efficiency and the accuracy of automated fact-checkers: The key breakthrough in effectively evaluating the factual accuracy of LLMs lies in establishing accurate and efficient fact-checkers. This requires improvement of the quality of the evidence used for making veracity decisions. Moreover, many recent methods rely on the factuality of stronger models such as GPT-4 for claim verification. Not only is this computationally expensive, but it also tends to be highly sensitive to minor prompt changes and LLM updates. A small task-specific and well fine-tuned NLI model can be a more viable, robust, and cost-efficient option.
@@ 7 Conclusion
We presented an overview on the factuality of LLMs, surveying a number of studies covering topics such as evaluation and improvement methods (applicable at various stages: pre-training, SFT, inference and post-processing) along with their respective challenges. We also identified three major issues and pointed out to promising future research directions.
Limitations
Despite conducting an extensive literature review to encompass all existing research on the factuality of LLMs, some studies may have been omitted due to the rapidly evolving nature of this research area. We endeavored to include all pertinent studies and references wherever feasible. This survey only briefly touches upon the factuality issues associated with vision language models. However, there is room for a more in-depth exploration of mitigation techniques specific to vision-language models. Additionally, comprehensive discussions are also necessary for language models that incorporate other modalities, such as video and speech.